OpenStudy (anonymous):

I need help with a word problem. During the first part of a trip a canoeist travels 35 miles at a certain speed. The canoeist travels 8miles on the second part of the trip at a speed 5mph slower. The total time of the trip is 4 hours. What was the speed of each trip?

OpenStudy (anonymous):

35/a+8/(a-5)=4

OpenStudy (anonymous):

m/(m/h)= h

OpenStudy (anonymous):

can you get it from there?

OpenStudy (anonymous):

I have no idea what to do. I don't understand what it means to do. I'm sorry I am such an idiot.

OpenStudy (anonymous):

ok first get a common denominator. to do this, the first equation (35/a) will have to be multiplied by (a-5)/(a-5)

OpenStudy (anonymous):

next the second part (8/(a-5)) must be multiplied by a/a

OpenStudy (anonymous):

this will allow you to add the terms together and do the next step. (35a-175+8a)/(a(a-5)) this simplifies to (43a-175)/(a^2-5a)=4

OpenStudy (anonymous):

then multiply each side by (a^2-5a) 43a-175=4(a^2-5a)

OpenStudy (anonymous):

43a-175=4a^2-20a then get it all on one side of the =. 4a^2-63a+175=0

OpenStudy (anonymous):

now all you have to do is use the quadratic form. and figure out which answer works.

OpenStudy (anonymous):

Thank you I think I have it from here. I hope any way.

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

t=d/t The total time is 4 hours. Solve the following for r, "a certain speed" for the 35 mile trip leg.\[\frac{35}{r}+\frac{8}{r-5}= 4\]Note: The answer for r is not a simple integer or fraction.\[r=\frac{1}{8} \left(63+\sqrt{1169}\right)=12.15 \text{ mph} \]\[r-5= 7.15 \text{ mph} \]

OpenStudy (anonymous):

\[\frac{35}{12.15}+\frac{8}{12.15-5}=4.000 \text{ hours} \text{ rounded} \text{ the} \text{ nearest } 1/1000 \]